OpenStudy (anonymous):

Use synthetic division to determine whether the number k is an upper or lower bound (as specified for the real zeros of the function f). k = 2; f(x) = 2x3 + 4x2 + 2x - 4; Lower bound? I believe that it is lower bound, is that correct? How would I solve this?

OpenStudy (mertsj):

Possible zeros: \[\pm\frac{4}{1},\pm\frac{4}{2},\pm\frac{2}{1},\pm\frac{2}{2}, \pm\frac{1}{1},\pm\frac{1}{2}\] or: \[\pm4, \pm2, \pm1, \pm\frac{1}{2}\]

OpenStudy (mertsj):

\[f(x)=2(x^2+2x^2+x-2)=0\] \[x^3+2x^2+x-2=0\]

OpenStudy (mertsj):



OpenStudy (mertsj):